ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 35, 46, 47, 58 Skrypnik Igor V., Institute of Applied Mathematics and Mechanics NASU, Donetsk, Ukraine Degree of densely defined (S)_{+} operators and applications. The introduction of the degree for densely defined operators of the class (S)_{+}, the calculation of the index of the critical points are given. We consider unbounded nonlinear operators with special unbounded linearization that is used for the evaluation of the index. The applications for nonlinear elliptic operators with strong coefficient growth and for the bifurcation problems are provided.